# A Linear Quantum Coupler for Clean Bosonic Control

### v2

@@ I Introduction
Driven superconducting circuits provide a platform for studying complex nonlinear quantum optics and manipulating quantum information in the microwave domain. The intrinsic nonlinearity of Josephson junctions [1] has been utilized to design numerous circuit elements, ranging from quantum-limited amplifiers [2, 3, 4, 5, 6, 7, 8, 9], novel non-reciprocal devices [10, 11, 12, 13], ultra-low-noise detectors [14, 15, 16], and, most notably, quantum information processing through the field of circuit Quantum Electrodynamics [17, 18]. A significant portion of these applications are enabled by nonlinear quantum mixers that activate specific desired processes when parametrically driven with a microwave drive. Such parametric mixers have found use as couplers [19, 20, 21, 22], amplifiers [9], and even nonlinear oscillators in their own right [23, 24, 25].
@@
A common approach to mitigate some parasitic effects is to use a ‘Kerr-free three wave-mixer’ [27, 28, 7, 29, 30], where one utilizes a third-order nonlinearity to generate the desired process with a single drive tone, while nulling the fourth-order (Kerr) nonlinearity. Suppressing the Kerr suppresses spurious processes like the AC Stark shift, which can lead to reduced saturation power in the amplifier context [31, 32], or frequency collisions with other sensitive modes in the coupler context [21]. However, for high-fidelity applications, this may still be insufficient – in such fragile systems, the loss of even a single photon to some uncontrolled degree of freedom can ruin sensitive quantum information. In particular, at strong drive powers, even Kerr-free mixers are inevitably spoiled by the onset of numerous transitions induced by nonlinearities beyond the fourth order. The effect of these transitions is especially evident at drive powers where the mixer ionizes [33, 34, 35, 26, 36] into higher-lying states or enters a chaotic regime, irreversibly spoiling the system’s information.
Figure 1: The LINC, a quantum single-balanced mixer a, Fundamentals of mixer balancing. By taking two identical nonlinear elements and driving them $180^{\circ}$ out of phase (e.g., with a balun), we can create a single-balanced mixer. If the nonlinear element is a single junction, this circuit is the familiar differentially-driven SQUID (DDS, [21]). If the nonlinear element is instead a DC flux-biased RF-SQUID, this circuit results in the Linear Inductive Coupler (LINC). b, An illustration of the advantage of a balanced mixer. In general, a driven nonlinear element will permit numerous nonlinear processes, of which only a small subset is typically desired. Balancing the nonlinear element will suppress a significant fraction of spurious processes (light orange), but preserve the desired process (black) at high strength. Any remaining parasitic processes (red) still permitted by the symmetry must be avoided through other means, like careful selection of drive frequencies. c, The LINC circuit, with a symmetric loop and capacitive pads, enables a separation of driven and undriven mixer behavior. At the operating point where the total DC flux in the outer SQUID loop is half a flux quantum, the loop junctions are biased to effectively infinite inductance, leaving only the linear shunt as the coupler’s inductor. This simultaneously nulls all nonlinearity when idle, which is beyond just being Kerr-free. When threaded with AC flux, the outer loop activates and drives a balanced three-wave mixing process.
@@ II The Linear Inductive Coupler
When the LINC is driven, the drive exactly cancels across the shunting inductor by design and does not displace the common-mode phase. The LINC’s Hamiltonian is thus separable into two non-interacting parts, the undriven and the driven, as shown in eq. 1. The driven portion contains beneficial symmetries due to the balancing of the drive, enforcing a selection rule on the allowed nonlinear processes in a manner similar to the DDS. Only processes where the number of coupler and resonator photons are even can be activated. Even within these, processes that are even-order in the drive, like the drive-induced frequency shift, are suppressed. This is in stark contrast to prevalent mixers like the SNAIL, where nominally every process can occur, but individual nonlinearities can be manually tuned to zero at specific flux bias points. We will show that this selection rule significantly expands the frequencies and strengths at which the LINC can be driven without causing spurious transitions, and reduce inter-modulation products in the presence of multiple drive tones. At leading order, the LINC functions as a three-wave mixer, activating either a beamsplitting (red sideband/ conversion) or squeezing (blue sideband/ gain) process when driven with a single tone of appropriate frequency.
We note that while the simultaneous balancing of the drive and the bias of this circuit is novel and crucial to its function, the circuit topology itself is fairly common in literature [20, 37, 38, 39, 40]. The LINC is a sibling to the Asymmetrically Threaded SQUID (ATS), with nominally identical circuit geometries but an exactly opposing notion of bias point. This allows the LINC to be optimized as a three-wave mixing coupler, while the ATS is utilized for Kerr-free four-wave mixing applications, like parametric two-photon dissipation [20]. The LINC also shares its topology with the gradiometric SNAIL [37], where the flux bias in the SQUID loops is used to tune the effective junction ratio of a regular charge-driven SNAIL, and symmetry is not a strict constraint. One key physical feature where the LINC might differ from these circuits is in its shunting inductor, which for the SNAIL and ATS typically consists of an array of a few junctions. Since bosonic applications might require highly linear idle circuits, the LINC shunt could benefit from using high kinetic inductance materials such as NbN or granular Aluminum (grAl), or a meandering geometric inductor, to minimize its static nonlinearity.
@@ III Idle nonlinearity and decoherence
Figure 2: Static Hamiltonian and flux noise sensitivity. a, The static LINC frequency and Kerr as a function of DC flux. Numerical values (solid lines) are calculated by exact diagonalization of a LINC Hamiltonian with $E_{L}/h=52.8$ GHz, $E_{J}/h=15.84$ GHz, $E_{C}/h=100$ MHz, with the center shunt composed of an array of 10 junctions, where $h$ is Planck’s constant. Overlayed analytic curves follow eqs. 3 and 4. At the operating point of $\phi_{\text{DC}}=\pi/2$, the coupler is linear. b, Three-wave mixing strength (dashed black) and sensitivity to flux-noise (yellow) as a function of DC flux. The former is calculated from a Floquet simulation of a parametric squeezing process. The latter is a simple derivative of the static frequencies calculated in part a ($1/2\pi\;d\omega_{L}/d\Phi$). These quantities are equal, up to a scaling factor of $2$ (see appendix A). c, Inherited dephasing for a coupled quantum memory, as a function of DC flux. Thermal noise-induced dephasing is calculated for a coupler thermal population of $2\%$ and $T_{1}$ of 20 $\mu$s, and is minimized at the operating point, where the coupler is linear. The low-frequency dephasing due to inherited flux noise is calculated through eq. 7, with noise amplitude $A_{\Phi}=1\mu\Phi_{0}/\sqrt{Hz}$. The latter dominates memory coherence, but may be suppressed through dynamical decoupling.
@@
[…]
@@
[…]
@@ IV The LINC as a driven mixer
Figure 3: The LINC as a driven mixer a, LINC beamsplitting strength ($g_{BS}$) and Kerr ($\alpha_{L}$) as a function of drive strength ($\phi_{\text{AC}}$), from an exact Floquet simulation. While the LINC is Kerr-free when idle, higher-order nonlinearities can induce a driven Kerr. This can be suppressed by arraying multiple LINCs while preserving the (driven) flux through each LINC loop (arrayed 3x, dashed). b, Driven LINC frequency shift (solid purple) vs drive strength. Unlike in charge-driven mixers like the SNAIL, this frequency shift is independent of the coupler Kerr, and persists even if the coupler is arrayed (dashed purple). It is possible to minimize this shift by biasing to a flux point where the coupler frequency has an inflection point as a function of DC flux (dotted black). c, Steady-state driven purity of the LINC (brown) and SNAIL (yellow) at constant drive amplitude ($\phi_{AC}=0.2\pi$) as a function of drive frequency. Both couplers are operated at a bias point where they are Kerr-free, and their parameters are optimized to match their beamsplitting strength and frequency at this bias point. Each coupler is simulated with equal decay, dephasing, and coupling to an external qubit ($\omega_{q}=4.9$ GHz). The LINC displays a significantly cleaner driven spectrum due to its parity protection. d, Steady-state driven purity of the LINC (brown) and SNAIL (yellow) in the presence of two drive tones (each with $\phi_{AC}=0.1\pi$), displaying the effect of parasitic intermodulation products. The two drive frequencies are scaled simultaneously maintaining a fixed ratio of 2:3, to ease Floquet-Markov simulations.
@@
where $\omega_{L}=\sqrt{8E_{L}E_{C}}$ is the frequency of the LINC mode at the operating point, $\hat{c}$ and $\hat{c}^{\dagger}$ are its ladder operators, $J_{1}(x)$ is the first-order Bessel function, and $\phi_{\text{AC}}(t)=|\phi_{\text{AC}}|\cos{\omega_{d}t}$ is the drive. For two modes Alice ($\hat{a}$, $\omega_{a}$) and Bob ($\hat{b}$, $\omega_{b}$) that are coupled to the LINC (with energy participations $p_{a}$ and $p_{b}$), driving at select frequencies can activate various desired processes, such as:
@@
Unfortunately, even an ideal three wave mixer incurs drive-induced frequency shifts, as can be seen from the average frequency of the above element $\langle\omega_{L}^{(\infty)}\rangle\propto\langle\sqrt{\left(1+2E_{J}/E_{L}\phi_{\text{AC}}\right)}\rangle\neq 0$. In fact, in contrast to a charge-driven element like the SNAIL, the driven frequency shift of the LINC is unaffected by arraying (see LABEL:fig3b), and is not related to the coupler’s anharmonicity. It is instead given by the susceptibility of the LINC frequency to change in the flux, $\Delta\omega_{L}\propto\left.d^{2}\omega/d\phi^{2}\right|_{\frac{\pi}{2}}$. One can minimize this driven shift (for example, when using the LINC in a parametric amplifier) by finding the inflection point $d^{2}\omega/d\phi^{2}\left[\phi_{\text{DC}}\right]=0$, which for the perfectly symmetric LINC (with $E_{J}/E_{L}=0.3$) occurs at $\phi_{\text{DC}}\sim 0.594\pi$ (dotted line in LABEL:fig3b). However, we note that this operating point is not parity-protected, and therefore might incur other parasitic processes that may be less desirable than the coupler frequency shift.
@@ V Effect of asymmetry and parasitic inductance
to lowest order in $\beta_{\Delta}$ and $\phi_{\Delta}$. This new minima results in a change in the undriven potential, meaning that Taylor expansions to compute parameters of interest should be expanded about this new point. In light of this shift, the static Kerr at the operating point becomes (see appendix C)
@@
for small asymmetries. Thus any static nonlinearity gained by the LINC in the presence of these asymmetries is only second-order in $\{\beta_{\Delta},\phi_{\Delta}\}$. Notably, the sign of the individual asymmetries is important, as the two effects can either constructively or destructively interfere. Additionally, in the absence of junction asymmetry, the static Kerr is nulled for any value of $\phi_{\Delta}$. As a reference, for a $\beta_{\Delta}=2\%$ and a $\phi_{\Delta}=-\frac{\pi}{2}\times 1\%$, the static Kerr of the LINC is roughly $-130$ KHz for an $E_{c}=100$ MHz. Additionally, the shifted Kerr-free point can be found by changing the symmetric flux bias by less than $0.005\pi$ (see appendix C).
When a LINC with non-zero asymmetry is driven, the LINC may turn on processes that would have otherwise been parity-protected. As an example, we consider the third-order mixing process involving two drive photons and a single LINC photon, corresponding to a parasitic sub-harmonic displacement of the coupler. This process is always allowed in any charge-driven mixer (like the SNAIL), yet is forbidden in the ideal LINC. The strength of this process, which we represent as $H_{\text{sub}}=g_{1,2}\phi_{\text{AC}}^{2}(\hat{c}+\hat{c}^{\dagger})$, is given by (appendix C):
@@
and is thus linearly sensitive to both junction and flux asymmetry, with their relative sensitivity depending on $\beta_{\Sigma}$. For modest asymmetries (again $\beta_{\Delta}=2\%$ and $\phi_{\Delta}=-\frac{\pi}{2}\times 1\%$), the relative suppression of such parasitic resonances is $g_{12}\phi_{\text{AC}}^{2}/g_{21}\phi_{\text{AC}}\approx 5\%$ for $\phi_{\text{AC}}=0.2\pi$, still achieving over an order of magnitude reduction compared to an ordinary 3-wave mixer.
Finally, for reasonable coupler geometries, one may expect the LINC to have a non-negligible parasitic linear inductance in series with the coupler. Such a parasitic inductor often has an impact beyond diluting the driven nonlinearity of a general coupler – for example, in the SNAIL, this can cause the Kerr-free operating point to shift significantly. In the LINC, this linear inductance is far less harmful, since it does not directly interact with the flux (drive) degree of freedom. In fact, any parasitic series inductance fully preserves the linearity of the idle LINC at the same operating point of $\phi_{\text{DC}}=\pi/2$, where every even order nonlinearity (including Kerr) is simultaneously suppressed. On the other hand, parasitic inductances within the LINC loop, while relatively small in magnitude, may shift the linear operating point slightly. We include a detailed analysis in appendices D and E. Overall, our simulations predict that the LINC remains a robust and protected quantum mixer in the presence of realistic experimental imperfections.

### v3

@@ I Introduction
Driven superconducting circuits provide a platform for studying complex nonlinear quantum optics and manipulating quantum information in the microwave domain. The intrinsic nonlinearity of Josephson junctions [1] has been utilized to design numerous circuit elements, ranging from quantum-limited amplifiers [2, 3, 4, 5, 6, 7, 8, 9], novel non-reciprocal devices [10, 11, 12, 13], ultra-low-noise detectors [14, 15, 16], and, most notably, quantum information processing through the field of circuit Quantum Electrodynamics (circuit QED) [17, 18]. A significant portion of these applications are enabled by nonlinear quantum mixers that activate specific desired processes when parametrically driven with a microwave drive. Such parametric mixers have found use as couplers [19, 20, 21, 22], amplifiers [9], and even nonlinear oscillators in their own right [23, 24, 25].
@@
A common approach to mitigate some parasitic effects is to use a ‘Kerr-free three-wave mixer’ [27, 28, 7, 29, 30], where one utilizes a third-order nonlinearity to generate the desired process with a single drive tone, while nulling the fourth-order (Kerr) nonlinearity. Suppressing the Kerr suppresses spurious processes like the AC Stark shift, which can lead to reduced saturation power in the amplifier context [31, 32], or frequency collisions with other sensitive modes in the coupler context [21]. However, for high-fidelity applications, this may still be insufficient – in such fragile systems, the loss of even a single photon to some uncontrolled degree of freedom can ruin sensitive quantum information. In particular, at strong drive powers, even Kerr-free mixers are inevitably spoiled by the onset of numerous transitions induced by nonlinearities beyond the fourth order. The effect of these transitions is especially evident at drive powers where the mixer ionizes [33, 34, 35, 26, 36] into higher-lying states or enters a chaotic regime, irreversibly spoiling the system’s information.
Figure 1: The LINC, a quantum single-balanced mixer a, Fundamentals of mixer balancing. By taking two identical nonlinear elements and driving them $180^{\circ}$ out of phase (e.g., with a balun), we can create a single-balanced mixer. If the nonlinear element is a single junction, this circuit is the familiar differentially-driven SQUID (DDS, [21]). If the nonlinear element is instead a DC flux-biased RF-SQUID, this circuit results in the Linear Inductive Coupler (LINC). The $\Pi$-T transformation provides a simpler circuit representation. b, An illustration of the advantage of a balanced mixer. In general, a driven nonlinear element will permit numerous nonlinear processes, of which only a small subset is typically desired. Balancing the nonlinear element will suppress a significant fraction of spurious processes (light orange), but preserve the desired process (black) at high strength. Any remaining parasitic processes (red) still permitted by the symmetry must be avoided through other means, like careful selection of drive frequencies. c, The LINC circuit, with a symmetric loop and capacitive pads, enables a separation of driven and undriven mixer behavior. At the operating point where the total DC flux in the outer SQUID loop is half a flux quantum, the loop junctions are biased to effectively infinite inductance, leaving only the linear shunt as the coupler’s inductor. This simultaneously nulls all nonlinearity when idle, which is beyond just being Kerr-free. When threaded with AC flux, the outer loop activates and drives a balanced three-wave mixing process.
@@ II The Linear Inductive Coupler
When the LINC is driven, the drive exactly cancels across the shunting inductor by design and does not displace the common-mode phase. The LINC’s Hamiltonian is thus separable into two non-interacting parts, the undriven and the driven, as shown in eq. 1. The driven portion contains beneficial symmetries due to the balancing of the drive, enforcing a selection rule on the allowed nonlinear processes in a manner similar to the DDS. Only processes where the number of coupler and resonator photons are even can be activated. Even within these, processes that are even-order in the drive, like the drive-induced frequency shift, are suppressed. This is in stark contrast to prevalent mixers like the SNAIL, where nominally every process can occur, but individual nonlinearities can be manually tuned to zero at specific flux bias points. We will show that this selection rule significantly expands the frequencies and strengths at which the LINC can be driven without causing spurious transitions, and reduce inter-modulation products in the presence of multiple drive tones. At leading order, the LINC functions as a three-wave mixer, activating either a beamsplitting (red sideband or conversion) or squeezing (blue sideband or gain) process when driven with a single tone of appropriate frequency.
We note that while the simultaneous balancing of the drive and the bias of this circuit is novel and crucial to its function, the circuit topology itself is fairly common in literature [20, 37, 38, 39, 40]. The LINC is a sibling to the Asymmetrically Threaded SQUID (ATS), with nominally identical circuit geometries but an exactly opposing notion of bias point. This allows the LINC to be optimized as a three-wave mixing coupler, while the ATS is utilized for Kerr-free four-wave mixing applications, like parametric two-photon dissipation [20].
The LINC also shares its topology with the gradiometric SNAIL [37], where the flux bias in the SQUID loops is used to tune the effective junction ratio of a regular charge-driven SNAIL, and symmetry is not a strict constraint. One key physical feature where the LINC might differ from these circuits is in its shunting inductor, which for the SNAIL and ATS typically consists of an array of a few junctions. While different circuits in the past have utilized inductive shunts for various improvements, here the shunt does not simply dilute nonlinearity. Due to the balanced nature of this circuit, this shunt entirely sets the linearity of its undriven Hamiltonian, and allows a suppression of unwanted interactions without reducing the strength of the desired process. Since bosonic applications might require highly linear idle circuits, the LINC shunt could benefit from using high kinetic inductance materials such as NbN or granular Aluminum (grAl), or a meandering geometric inductor, to minimize its static nonlinearity.
@@ III Idle nonlinearity and decoherence
[…]
@@
$\begin{split}1-\mathcal{F}_{p}&\leq\left.\left(\dfrac{2\pi}{g_{\text{3wm}}}\dfrac{dg_{\text{3wm}}}{d\Phi}\right)^{2}\right|_{\frac{\pi}{2}}\int_{0}^{\infty}S_{\Phi\Phi}[f]\;g_{N}[f\tau_{\text{g}}]\>df\\ &\approx\left(\dfrac{2E_{J}}{E_{L}}\right)^{2}\int_{1/\tau_{\text{exp}}}^{1/\tau_{\text{g}}}\frac{1}{\Phi_{0}^{2}}\;S_{\Phi\Phi}[f]\>df.\end{split}$ (6)
@@
$\begin{split}\kappa_{\varphi}&\propto\sqrt{\left(d\omega_{\text{res}}/d\Phi\right)^{2}\int_{0}^{\infty}S_{\Phi\Phi}[f]\;g_{N}[f\tau]\>df}\\ &\approx p_{\text{res}}\left|\frac{d\omega_{L}}{d\Phi}\right|_{\phi_{\text{DC}}=\pi/2}A_{\Phi}\;C,\end{split}$ (7)
@@ IV The LINC as a driven mixer
Figure 3: The LINC as a driven mixer a, LINC beamsplitting strength ($g_{BS}$) and Kerr ($\alpha_{L}$) as a function of drive strength ($\phi_{\text{AC}}$), from an exact Floquet simulation. While the LINC is Kerr-free when idle, higher-order nonlinearities can induce a driven Kerr. This can be suppressed by arraying multiple LINCs while preserving the (driven) flux through each LINC loop (arrayed 3x, dashed). b, Driven LINC frequency shift (solid purple) vs drive strength. Unlike in charge-driven mixers like the SNAIL, this frequency shift is independent of the coupler Kerr, and persists even if the coupler is arrayed (dashed purple). It is possible to minimize this shift by biasing to a flux point where the coupler frequency has an inflection point as a function of DC flux (dotted black). c, Driven steady-state impurity of the LINC (left) and an equivalent SNAIL (right) as a function of drive frequency and amplitude. Both couplers are operated at a bias point where they are Kerr-free, and their parameters are optimized to match their beamsplitting strength and frequency at this bias point. Each coupler is simulated with equal decay and low-frequency dephasing. The LINC displays a significantly cleaner driven spectrum due to its parity protection.
@@
where $\omega_{L}=\sqrt{8E_{L}E_{C}}$ is the frequency of the LINC mode at the operating point and $\hat{c}$ and $\hat{c}^{\dagger}$ are its ladder operators. $J_{1}(x)$ is the first-order Bessel function arising from the sinusoidal dependence of the drive in the LINC Hamiltonian, with $\phi_{\text{AC}}(t)=|\phi_{\text{AC}}|\cos{\omega_{d}t}$ being the drive. For two modes Alice ($\hat{a}$, $\omega_{a}$) and Bob ($\hat{b}$, $\omega_{b}$) that are coupled to the LINC (with energy participations $p_{a}$ and $p_{b}$), driving at select frequencies can activate various desired processes, such as:
@@
Unfortunately, even an ideal three wave mixer incurs drive-induced frequency shifts, as can be seen from the average frequency of the above element $\langle\omega_{L}^{(\infty)}\rangle\propto\langle\sqrt{\left(1+2E_{J}/E_{L}\phi_{\text{AC}}\right)}\rangle\neq 0$. In fact, in contrast to a charge-driven element like the SNAIL, the driven frequency shift of the LINC is unaffected by arraying (see LABEL:fig3b), and is not related to the coupler’s anharmonicity. It is instead given by the susceptibility of the LINC frequency to change in the flux, $\Delta\omega_{L}\propto\left.d^{2}\omega/d\phi^{2}\right|_{\frac{\pi}{2}}$. One can minimize this driven shift (for example, when using the LINC in a parametric amplifier) by finding the inflection point $d^{2}\omega/d\phi^{2}\left[\phi_{\text{DC}}\right]=0$, which for the perfectly symmetric LINC (with $E_{J}/E_{L}=0.3$) occurs at $\phi_{\text{DC}}\sim 0.594\pi$ (dotted line in LABEL:fig3b). However, we note that this operating point is not Kerr-free and it breaks the weaker parity protection with respect to the drive, and therefore might suffer parasitic processes less desirable than the coupler frequency shift.
@@ V Effect of asymmetry and parasitic inductance
to lowest order in $\beta_{\Delta}$ and $\phi_{\Delta}$. This new minima results in a change in the undriven potential, meaning that Taylor expansions to compute parameters of interest should be expanded about this new point. In light of this shift, the static Kerr at the operating point becomes (see appendix D)
@@
for small asymmetries. Thus any static nonlinearity gained by the LINC in the presence of these asymmetries is only second-order in $\{\beta_{\Delta},\phi_{\Delta}\}$. Notably, the sign of the individual asymmetries is important, as the two effects can either constructively or destructively interfere. Additionally, in the absence of junction asymmetry, the static Kerr is nulled for any value of $\phi_{\Delta}$. As a reference, for a $\beta_{\Delta}=2\%$ and a $\phi_{\Delta}=-\frac{\pi}{2}\times 1\%$, the static Kerr of the LINC is roughly 130-130 KHz for an $E_{c}=100$ MHz. Additionally, the shifted Kerr-free point can be found by changing the symmetric flux bias by less than $0.005\pi$ (see appendix D).
When a LINC with non-zero asymmetry is driven, the LINC may turn on processes that would have otherwise been parity-protected. As an example, we consider the third-order mixing process involving two drive photons and a single LINC photon, corresponding to a parasitic sub-harmonic displacement of the coupler. This process is always allowed in any charge-driven mixer (like the SNAIL), yet is forbidden in the ideal LINC. The strength of this process, which we represent as $H_{\text{sub}}=g_{1,2}\phi_{\text{AC}}^{2}(\hat{c}+\hat{c}^{\dagger})$, is given by (appendix D):
@@
and is thus linearly sensitive to both junction and flux asymmetry, with their relative sensitivity depending on $\beta_{\Sigma}$. For modest asymmetries (again $\beta_{\Delta}=2\%$ and $\phi_{\Delta}=-\frac{\pi}{2}\times 1\%$), the relative suppression of such parasitic resonances is $g_{12}\phi_{\text{AC}}^{2}/g_{21}\phi_{\text{AC}}\approx 5\%$ for $\phi_{\text{AC}}=0.2\pi$, still achieving over an order of magnitude reduction compared to an ordinary 3-wave mixer. In effect, since the LINC is a balanced RF-SQUID, any asymmetry only brings its performance closer to the performance of a regular RF-SQUID or SNAIL. The full effect of asymmetries on parasitic resonances can be captured by plotting the available drive-space for the LINC, similar to LABEL:appx_fig3, which add no significant spurious processes up to $\phi_{\text{AC}}=0.2\pi$ even at junction asymmetries up to $10\%$.
Finally, for reasonable coupler geometries, one may expect the LINC to have a non-negligible parasitic linear inductance in series with the coupler. Such a parasitic inductor often has an impact beyond diluting the driven nonlinearity of a general coupler – for example, in the SNAIL, this can cause the Kerr-free operating point to shift significantly. In the LINC, this linear inductance is far less harmful, since it does not directly interact with the flux (drive) degree of freedom. In fact, any parasitic series inductance fully preserves the linearity of the idle LINC at the same operating point of $\phi_{\text{DC}}=\pi/2$, where every even order nonlinearity (including Kerr) is simultaneously suppressed. On the other hand, parasitic inductances within the LINC loop, while relatively small in magnitude, may shift the linear operating point slightly. We include a detailed analysis in appendices E and F. Overall, our simulations predict that the LINC remains a robust and protected quantum mixer in the presence of realistic experimental imperfections.